# SMReferences for Appendix

### v3

@@ 4.1 Regression coefficient
As mentioned in the Introduction, our Double-Calibration strategy can be extended to target parameters other than ATE. In this section, we demonstrate the broad applicability of our approach via another important parameter in statistics and econometrics. Specifically, we consider the following semiparametric partially linear model (Robinson, 1988):
@@
$\hat{\tau}_{\textrm{DCal}}\coloneqq\hat{\tau}+\frac{1}{n}\sum_{i=1}^{n}\frac{(T_{i}-\tilde{\pi}_{i})(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\sigma}_{e}^{2}},$
where $\hat{\mu}$ is estimated by solving the same constrained quadratic program as in (5) for ATE, except that we replace the diagonal elements of ${\bm{\Pi}}$ with $\phi^{\prime}(X_{i}\hat{\gamma})$ and $\tilde{R}_{i}$ with $Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}$; $\tilde{\sigma}_{e}^{2}$ is an estimated variance using $\tilde{\pi}$, i.e.,
@@
Here we deliberately choose not to write $\tilde{r}$ in place of $\hat{r}+\hat{\mu}$ as the purpose of $\hat{\mu}$ is to correct the estimation error of $T\hat{\tau}+\hat{r}$ as an estimator of $\pi^{\ast}\tau^{\ast}+r^{\ast}$. This is slightly different from the setting of ATE in Section 2 and 3. ∎
We now turn to the nuisance function $\tilde{\pi}$, which is constructed by modifying the estimator in (8) slightly. Specifically, we keep the objective function and the first three constraints the same as before, except that the term $T_{i}/\pi_{i}-1$ becomes $T_{i}-\pi_{i}$. As for the last constraint (12), we break it into the following two constraints:
$\begin{split}&\max_{i}|\pi_{i}|\leq M_{\pi},\quad\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\pi_{i})^{2}\geq M_{\pi}^{-1},\\ &\left|\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\pi_{i})\pi_{i}\right|\leq\eta_{\pi_{1}}\cdot\frac{\|{\bm{\pi}}\|_{2}}{\sqrt{n}}.\end{split}$ (16)
To understand the theoretical property of the DCal estimator of the regression coefficient $\tau^{\ast}$, we need to further modify the assumptions in Section 3. In particular, we adapt Conditions 4 and 6 to the following:
Condition 10.
@@
    There exists a constant $\hat{m}>0$ such that with probability converging to $1$, $|\hat{\tau}|\leq\hat{m}$ and that for all $i$,
    $\max\{|\hat{r}_{i}|,|\hat{\pi}_{i}|,|\hat{r}_{\mathrm{aux},i}|,|\hat{\pi}_{\mathrm{aux},i}|\}\leq\hat{m}.$
With the above preparation, by mimicking the proofs of Theorems 3 and 5 (see Appendix A.6), we have the following results on the asymptotic statistical property of $\hat{\tau}_{\textrm{DCal}}$ for estimating the $\tau^{\ast}$ in the semiparametric partially linear model (15).
@@
Under Conditions 5 and 10, by choosing $\eta_{r},\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{r},M_{\pi}\asymp 1$ sufficiently large, we have the following:
@@
    If Condition 7 holds, with the additional assumption that $|\hat{\tau}-\tau^{*}|=O(\sqrt{s_{r}\log p/n})$, then we have
    $\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{(T_{i}-\tilde{\pi}_{i})\varepsilon_{i}}{\tilde{\sigma}_{e}^{2}}+o_{\mathbb{P}}(1).$
@@
    If Conditions 8 and 9 holds, then we have
    $\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}\varepsilon_{i}}{\sigma_{e}^{2}}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})}{\sigma_{e}^{2}}+o_{\mathbb{P}}(1).$
We first discuss the intuitions of Theorem 6(i). With a slight abuse of notation, we write $\bar{\sigma}_{r}^{2}\coloneqq\frac{\textrm{var}(\varepsilon)}{\hat{\sigma}_{\varepsilon}^{2}}$. Analogous to the discussions below Theorem 3, informally, it says that
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\tau^{*})\mid\{(\mathbf{T},\mathbf{X}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{r}^{2}),$
@@
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\tau^{*})\mid\{(\mathbf{T},\mathbf{X}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\approx\mathcal{N}(0,\bar{\sigma}_{\pi}^{2}+\sigma_{\mu}^{2}),$
@@
Since all the summands in $\sigma_{\mu}^{2}$ are bounded with probability converging to $1$, $\sigma_{\mu}^{2}$ is with probability converging to $1$ of order $O(1)$, which in turn implies that $\hat{\tau}_{\textrm{DCal}}$ is a $\sqrt{n}$-consistent estimator of $\tau^{*}$.
@@
Our result is also related to the vast literature on debiased lasso (Zhang and Zhang, 2014; van de Geer et al., 2014; Javanmard and Montanari, 2014, 2018; Cai and Guo, 2017; Zhu and Bradic, 2018; Cattaneo et al., 2018, 2019). We refer readers who are interested in a comparison between our result and the above references to Section 1 of Bradic et al. (2019a), which has comprehensively surveyed the conditions under which $\sqrt{n}$-consistent estimator of $\tau^{\ast}$ were developed in some of these referenced works. ∎
Remark 8.
Last but not least, we briefly comment on the computational issue of the above constrained program for calibrating regression coefficient estimation. Since the constraint (16) is non-convex, the time complexity of the resulted optimization problem is generally NP-hard. Note also that the above computational issue is also related to ATE estimation in Section 2 and 3. Depending on the link functions, the resulting optimization problem for Double-Calibration can also be non-convex. Smucler et al. (2019) also discussed extensively on the potential drawback of solving non-convex program, such as the lack of theoretical convergence guarantees. This is a very important and potentially difficult problem that does not catch enough attention in the causal inference literature. But in order not to lead readers astray from the main message of the paper, we leave this problem to future work. ∎
4.1.1 Further explanation on the DCal estimator for the regression coefficient
We expect that it may be helpful to explain the DCal estimator for regression coefficient estimation from the perspective of the influence function of DR functionals (Rotnitzky et al., 2021). Generically, the influence function of a DR functional $\tau$ with nuisance parameters $\theta=(a,b)$ has the following form:
@@ 4.1.1 Further explanation on the DCal estimator for the regression coefficient
where the observable is $O=(Z,W)$, and $a\mapsto m_{a}(O,a)$ and $b\mapsto m_{b}(O,b)$ are two linear maps that satisfy some extra regularity and moment conditions (which are to ensure $\mathsf{IF}$ is doubly robust). We refer readers to Rotnitzky et al. (2021) and Liu et al. (2023) for more details.
@@
$\hat{\tau}_{\textrm{DML}}\coloneqq\hat{\tau}+\frac{1}{n}\sum_{i=1}^{n}\hat{a}(Z_{i})\left(Y_{i}-\hat{b}(Z_{i})\right)$
where $\hat{b}(z)\equiv\hat{r}(x)+\hat{\tau}t\equiv\psi(x^{\top}\hat{\beta})+\hat{\tau}t$ is simply some variant of the lasso estimate with the coefficients $(\hat{\tau},\hat{\beta}^{\top})^{\top}$ computed from the training dataset $\mathcal{D}_{\mathrm{tr}}$. As for $\hat{a}(z)$, as is done in this paper, one can simply first estimate $\hat{\gamma}$ to obtain $\hat{\pi}$ and then estimate the denominator $\mathbb{E}[(T-\pi^{\ast}(X))^{2}]$ by its DML estimator $n^{-1}\sum_{i=1}^{n}(T_{i}-\phi(X_{i}^{\top}\hat{\gamma}))^{2}$. To achieve $\sqrt{n}$-rate under $s_{r}\wedge s_{\pi}\ll\sqrt{n}/\log p$, one uses the Double-Calibration program to bias-correcting both $\hat{a}$ and $\hat{b}$ by using $\tilde{a}(Z)=(T-\phi(X^{\top}\tilde{\gamma})))/n^{-1}\sum_{i=1}^{n}(T_{i}-\phi(X_{i}^{\top}\tilde{\gamma}))^{2}$ and $\tilde{b}=\hat{b}(Z)+\hat{\mu}$ instead. By design of the Double-Calibration program, $\tilde{a}$ and $\tilde{b}$ again satisfy the following moment equations, similar to those in Remark 1 and 2:

### v4

@@ 4.1 Regression coefficient
As mentioned in the Introduction, our DCal strategy can be extended to target parameters other than ATE. In this section, we demonstrate the broad applicability of our approach via another important parameter in statistics and econometrics. Specifically, we consider the following semiparametric partially linear model (Robinson, 1988):
@@
$\hat{\tau}_{\mathrm{DCal}}\coloneqq\hat{\tau}+\frac{1}{n}\sum_{i=1}^{n}\frac{(T_{i}-\tilde{\pi}_{i})(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})}{\tilde{\sigma}_{e}^{2}},$
where $\hat{\bm{\mu}}$ is the solution to the following $\ell_{\infty}$-calibration program:
$\hat{\bm{\mu}}\coloneqq\mathcal{L}_{\bm{\mu},B_{\infty}^{n}(M_{r})}\left(\|\cdot\|_{2}^{2};\Delta^{(\mathrm{aux})}_{1,Y-T_{i}-\hat{\tau}-\hat{r},\mu};\phi^{\prime}(X^{\top}\hat{\gamma})X;\mathbbm{1}_{p}\eta_{r}\right)$
and $\tilde{\sigma}_{e}^{2}$ is an estimated variance using $\tilde{\pi}$, i.e.,
@@
Here we deliberately choose not to write $\tilde{r}$ in place of $\hat{r}+\hat{\mu}$ as the purpose of $\hat{\mu}$ is to correct the estimation error of $T\hat{\tau}+\hat{r}$ as an estimator of $\pi^{\ast}\tau^{\ast}+r^{\ast}$. This is slightly different from the setting of ATE in Sections 2 and 3. ∎
We now turn to the nuisance function $\tilde{\pi}=\phi(\tilde{X}^{\top}\tilde{\gamma})$, which is constructed by modifying the estimator in (10) slightly as follows:
$\tilde{\gamma}\coloneqq\mathcal{L}_{\gamma,\tilde{B}_{\infty}^{{\dagger}p}(M_{\pi})}\left(\|\cdot-\hat{\gamma}\|_{1};\Delta_{1,T,\phi(X^{\top}\gamma)};\mathrm{h};\mathrm{b}\right)$ (14)
where
$\tilde{B}_{\infty}^{{\dagger}p}(M_{\pi})$$\coloneqq\left\{\gamma:\|\bm{\pi}\|_{\infty}\leq M_{\pi}\right\},$ (14a)
$\mathrm{h}$ $\coloneqq\left(\psi^{\prime}(X^{\top}\hat{\beta})X^{\top},\hat{\mu},\tilde{X}^{\top},\pi,\pi-T\right)^{\top},$ (14b)
$\mathrm{b}$ $\coloneqq\left(\frac{\eta_{\pi_{1}}}{\sqrt{n}}\|\psi^{\prime}(X^{\top}\hat{\beta})X\|_{n,2},\frac{\eta_{\pi_{1}}\|\hat{\bm{\mu}}\|_{2}}{\sqrt{n}},\frac{\eta_{\pi_{2}}}{\sqrt{n}}\|\tilde{X}\|_{n,2},\frac{\eta_{\pi_{1}}\|\hat{\bm{\pi}}\|_{2}}{\sqrt{n}},-M_{\pi}^{-1}\right)^{\top}.$ (14c)
Specifically, we keep the objective function  (10), change the residual to be calibrated from $\Delta_{2,T,\pi}=T/\pi-1$ to $\Delta_{1,T,\pi}=T-\pi$, and add two extra inequality constraints to $\mathrm{h}$ and $\mathrm{b}$. In particular, the additional constraints are:
$\left|\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\pi_{i})\pi_{i}\right|\leq\eta_{\pi_{1}}\cdot\frac{\|{\bm{\pi}}\|_{2}}{\sqrt{n}},\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\pi_{i})^{2}\geq M_{\pi}^{-1}.$ (16)
To understand the theoretical property of the DCal estimator of the regression coefficient $\tau^{\ast}$, we need to further modify the assumptions in Section 3. In particular, we adapt Conditions 4(i) and 5 to the following:
Condition 9.
@@
    There exists a constant $\hat{m}>0$ such that with probability converging to $1$, $|\hat{\tau}|\leq\hat{m}$, $\mathbb{P}(|\psi(X^{\top}\hat{\beta})|\leq\hat{m}\mid\hat{\beta})=1$ and that $\mathbb{P}(|\phi(X^{\top}\hat{\gamma})|\leq\hat{m}\mid\hat{\gamma})=1$, where $X$ follows the same interpretation as Condition 5.
With the above preparation, by mimicking the proofs of Theorems 3 and 5 (see Appendix B.1), we have the following results on the asymptotic statistical property of $\hat{\tau}_{\mathrm{DCal}}$ for estimating the $\tau^{\ast}$ in the semiparametric partially linear model (13).
@@
Under Conditions 4(ii) and 9, by choosing $\eta_{r},\eta_{\pi_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}},M_{r},M_{\pi}\asymp 1$ sufficiently large, we have the following:
@@
    If Condition 6 holds, with the additional assumption that $|\hat{\tau}-\tau^{*}|=O(\sqrt{s_{r}\log p/n})$ and that $s_{r}=o(\sqrt{n}/\log p)$, we have
    $\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{(T_{i}-\tilde{\pi}_{i})\varepsilon_{i}}{\tilde{\sigma}_{e}^{2}}+o_{\mathbb{P}}(1).$
@@
    If Conditions 7 and 8 holds, then we have
    $\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}\varepsilon_{i}}{\sigma_{e}^{2}}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})}{\sigma_{e}^{2}}+o_{\mathbb{P}}(1).$
We first discuss the intuitions of Theorem 6 (i). With a slight abuse of notation, we write $\bar{\sigma}_{r}^{2}\coloneqq\frac{\mathrm{var}(\varepsilon)}{\hat{\sigma}_{\varepsilon}^{2}}$. Analogous to the discussions below Theorem 3, informally, it says that
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\tau^{*})\mid\{(\mathbf{T},\mathbf{X}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\overset{\mathsf{d}}{\approx}\mathcal{N}(0,\bar{\sigma}_{r}^{2}),$
@@
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\tau^{*})\mid\{(\mathbf{T},\mathbf{X}),\mathcal{D}_{\mathrm{aux}},\mathcal{D}_{\mathrm{tr}}\}\overset{\mathsf{d}}{\approx}\mathcal{N}(0,\bar{\sigma}_{\pi}^{2}+\sigma_{\mu}^{2}),$
@@
Since all the summands in $\sigma_{\mu}^{2}$ are bounded with probability converging to $1$, $\sigma_{\mu}^{2}$ is with probability converging to $1$ of order $O(1)$, which in turn implies that $\hat{\tau}_{\mathrm{DCal}}$ is a $\sqrt{n}$-consistent estimator of $\tau^{*}$.
@@
Our result is also related to the vast literature on debiased lasso (Zhang and Zhang, 2014; van de Geer et al., 2014; Javanmard and Montanari, 2014, 2018; Cai and Guo, 2017; Zhu and Bradic, 2018; Cattaneo et al., 2018, 2019; Shah and Bühlmann, 2023). We refer readers who are interested in a comparison between our result and the above references to Section 1 of Bradic et al. (2019a), which has comprehensively surveyed the conditions under which $\sqrt{n}$-consistent estimators of $\tau^{\ast}$ were developed in some of these referenced works. ∎
Further explanation on the DCal estimator for the regression coefficient through the lens of DR functionals
We expect that it may be helpful to explain the DCal estimator for regression coefficient estimation through the lens of influence functions of DR functionals, first characterized in Rotnitzky et al. (2021) (also see Section 5 of Chernozhukov et al. (2022a)). The influence function of a DR functional $\tau$ with nuisance parameters $\theta=(a,b)$ has the following generic form:
@@
where the observable is $O=(Z,W)$, and $a\mapsto m_{a}(O,a)$ and $b\mapsto m_{b}(O,b)$ are two linear maps that satisfy some extra regularity and moment conditions. We refer readers to Rotnitzky et al. (2021) and Liu et al. (2024) for more details.
@@
$\hat{\tau}_{\mathrm{DML}}\coloneqq\hat{\tau}+\frac{1}{n}\sum_{i=1}^{n}\hat{a}(Z_{i})\left(Y_{i}-\hat{b}(Z_{i})\right)$
where $\hat{b}(z)\equiv\hat{r}(x)+\hat{\tau}t\equiv\psi(x^{\top}\hat{\beta})+\hat{\tau}t$ is simply some variant of the lasso estimate with the coefficients $(\hat{\tau},\hat{\beta}^{\top})^{\top}$ computed from the training dataset $\mathcal{D}_{\mathrm{tr}}$. As for $\hat{a}(z)$, as is done in this paper, one can simply first estimate $\hat{\gamma}$ to obtain $\hat{\pi}$ and then estimate the denominator $\mathbb{E}[(T-\pi^{\ast}(X))^{2}]$ by its DML estimator $n^{-1}\sum_{i=1}^{n}(T_{i}-\phi(X_{i}^{\top}\hat{\gamma}))^{2}$. To achieve $\sqrt{n}$-rate under $s_{r}\wedge s_{\pi}\ll\sqrt{n}/\log p$, one uses the Double-Calibration program to de-bias $\hat{a}(Z)$ and $\hat{b}(Z)$ by, respectively, $\tilde{a}(Z)=(T-\phi(X^{\top}\tilde{\gamma})))/n^{-1}\sum_{i=1}^{n}(T_{i}-\phi(X_{i}^{\top}\tilde{\gamma}))^{2}$ and $\tilde{b}(Z)=\hat{b}(Z)+\hat{\mu}$. By design of the Double-Calibration program, $\tilde{a}$ and $\tilde{b}$ again satisfy the following moment equations, similar to those in Remark 1 and 2:
@@
4.2 Local Average Treatment Effects (LATE) with Instrumental Variables (IV)
Now we turn our attention to the setting allowing endogeneity by leveraging the availability of an Instrumental Variable (IV). Here we have access to $N$ i.i.d. observations $\{O_{i}=(X_{i},Z_{i},T_{i},Y_{i}),i=1,\cdots,N\}$, where $T\in\{0,1\}$ is again the binary treatment and $Z\in\{0,1\}$ is a binary IV. We are interested in estimating the so-called Local Average Treatment Effects (LATE) $\chi^{*}:=\mathbb{E}[Y(1)\mid T(1)>T(0)]$ (Imbens and Angrist, 1994). Under some standard IV-identification conditions which we defer to the Appendix B.2, $\chi^{*}$ is identifiable in the form of “2-Stage Least Squares” (2SLS):
$\chi^{\ast}=\frac{\mathbb{E}[Y(1)(T(1)-T(0))]}{\mathbb{E}[(T(1)-T(0))]}\coloneqq\frac{\mathbb{E}\left[\frac{ZTY}{\zeta^{\ast}(X)}-\frac{(1-Z)TY}{1-\zeta^{\ast}(X)}\right]}{\mathbb{E}\left[\frac{ZT}{\zeta^{\ast}(X)}-\frac{(1-Z)T}{1-\zeta^{\ast}(X)}\right]},$ (17)
where $\zeta^{\ast}(x)\coloneqq\mathbb{E}[Z|X=x]\equiv\varphi(x^{\top}\xi^{\ast})$ represents the Instrument Propensity Score (IPS). It is worth noting that $\chi^{\ast}$ may identify ATE instead of LATE under certain alternative IV identification conditions (Wang and Tchetgen Tchetgen, 2018).
To ease exposition, we consider the following parameter instead, which shares exactly the same structure as LATE:
$\tau^{*}\coloneqq\frac{\mathbb{E}[Y(1)T(1)]}{\mathbb{E}[T(1)]}\equiv\frac{\mathbb{E}\left[\frac{ZTY}{\zeta^{\ast}(X)}\right]}{\mathbb{E}\left[\frac{ZT}{\zeta^{\ast}(X)}\right]}\eqqcolon\frac{\tau_{n}^{*}}{\tau_{d}^{*}}.$
We next introduce the following notations for nuisance parameters involved in estimating $\tau^{\ast}$:
$\pi^{\ast}(x)$ $\coloneqq\mathbb{E}[T|X=x,Z=1]\equiv\phi(x^{\top}\gamma^{\ast}),$
$r^{\ast}(x)$ $\coloneqq\mathbb{E}[Y|X=x,Z=1,T=1]\equiv\psi(x^{\top}\beta^{\ast}).$
Then applying again the standard IV-identification conditions, we have
$\tau^{\ast}_{n}\equiv\mathbb{E}\left[\frac{Z(TY-\pi^{\ast}(X)r^{\ast}(X))}{\zeta^{\ast}(X)}+\pi^{\ast}(X)r^{\ast}(X)\right]\quad\&\quad\tau^{\ast}_{d}\equiv\mathbb{E}\left[\frac{Z(T-\pi^{\ast}(X))}{\zeta^{\ast}(X)}+\pi^{\ast}(X)\right].$
This gives rise to the following DML-2SLS estimator of $\tau^{\ast}$ via estimating $\tau_{n}^{\ast}$ and $\tau_{d}^{\ast}$ in (17) through their respective DML estimators (Tan, 2006): $\hat{\tau}_{\mathrm{DML}}\coloneqq\hat{\tau}_{n,\mathrm{DML}}/\hat{\tau}_{d,\mathrm{DML}}$, where
$\hat{\tau}_{n,\mathrm{DML}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{Z_{i}}{\hat{\zeta}_{i}}\left(T_{i}Y_{i}-\hat{\pi}_{i}\hat{r}_{i}\right)+\hat{\pi}_{i}\hat{r}_{i},\quad\&\quad\hat{\tau}_{d,\mathrm{DML}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{Z_{i}}{\hat{\zeta}_{i}}\left(T_{i}-\hat{\pi}_{i}\right)+\hat{\pi}_{i}.$
Following standard arguments in doubly robust estimation (Chernozhukov et al., 2018; Liu et al., 2024), it is straightforward to see that $\hat{\tau}_{\mathrm{DML}}$ is $\sqrt{n}$-consistent if $s_{\zeta}\cdot(s_{\pi}+s_{r})=o(n/\log^{2}p)$. Similar to ATE, we are interested in whether it is possible to construct $\sqrt{n}$-consistent estimators for LATE, under the laxed sparsity condition $s_{\zeta}\wedge(s_{\pi}\vee s_{r})=o(\sqrt{n}/\log p)$. Indeed, following similar arguments to those in Section 2, we consider the following Double-Calibrated estimator for LATE: $\hat{\tau}_{\textrm{DCal}}\coloneqq\hat{\tau}_{n,\textrm{DCal}}/\hat{\tau}_{d,\textrm{DCal}}$, where
$\hat{\tau}_{n,\textrm{DCal}}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{Z_{i}}{\tilde{\zeta}_{i}}\left(T_{i}Y_{i}-\hat{\pi}_{i}\hat{r}_{i}-\tilde{\mu}_{ty,i}\right)+\hat{\pi}_{i}\hat{r}_{i}+\tilde{\mu}_{ty,i},$
$\hat{\tau}_{d,1,\textrm{DCal}}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\frac{Z_{i}}{\tilde{\zeta}_{i}}\left(T_{i}-\hat{\pi}_{i}-\tilde{\mu}_{t,i}\right)+\hat{\pi}_{i}+\tilde{\mu}_{t,i}.$
Here, for $i=1,\cdots,n$, $\bm{\tilde{\mu}}_{ty}\coloneqq(\tilde{\mu}_{ty,1},\cdots,\tilde{\mu}_{ty,n})^{\top}$ and $\bm{\tilde{\mu}}_{t}\coloneqq(\tilde{\mu}_{t,1},\cdots,\tilde{\mu}_{t,n})^{\top}$ are, respectively, the solution to:
$\tilde{\bm{\mu}}_{ty}$ $\coloneqq\mathcal{L}_{\bm{\mu},B_{\infty}^{n}(M_{r})}\left(\|\cdot\|_{2}^{2};\Delta_{1,Z(TY-\hat{\pi}\hat{r})/\hat{\zeta},\mu_{ty}}^{(\mathrm{aux})};\varphi^{\prime}(X^{\top}\hat{\xi})X/\hat{\zeta};\mathbbm{1}_{p}^{\top}\eta_{ty}\right),$
$\tilde{\bm{\mu}}_{t}$ $\coloneqq\mathcal{L}_{\bm{\mu},B_{\infty}^{n}(M_{\pi})}\left(\|\cdot\|_{2}^{2};\Delta_{1,Z(T-\hat{\pi})/\hat{\zeta},\mu_{t}}^{(\mathrm{aux})};\varphi^{\prime}(X^{\top}\hat{\xi})X/\hat{\zeta};\mathbbm{1}_{p}^{\top}\eta_{t}\right).$
Similarly, $\tilde{\zeta}_{i}\coloneqq\varphi(\tilde{\xi}^{\top}\tilde{X}_{i})$, with $\tilde{\xi}$ as the solution to:
$\tilde{\xi}\coloneqq\mathcal{L}_{\xi,\tilde{B}_{\infty}^{p}(M_{\zeta})}\left(\|\cdot-\hat{\xi}\|_{1};\Delta_{2,Z,\varphi(X^{\top}\xi)};\mathrm{h};\mathrm{b}\right),$
where
$\tilde{B}_{\infty}^{p}(M_{\zeta})$$\coloneqq\left\{\gamma:\|Z/\varphi(\tilde{X}^{\top}\xi)\|_{\infty}\leq M_{\zeta}\right\},$
$\mathrm{h}$ $\coloneqq\left(\phi^{\prime}(X^{\top}\hat{\gamma})X^{\top},\psi^{\prime}(X^{\top}\hat{\beta})\hat{\pi}X^{\top},\phi^{\prime}(X^{\top}\hat{\gamma})\hat{r}X^{\top},\tilde{\mu}_{t},\tilde{\mu}_{ty},\tilde{X}^{\top}\right)^{\top},$
and $\text{and }\mathrm{b}$ $\coloneqq\left(\frac{\eta_{\zeta_{1}}}{\sqrt{n}}\|\psi^{\prime}(X^{\top}\hat{\beta})X\|_{n,2},\frac{\eta_{\zeta_{1}}\|\tilde{\bm{\mu}}_{t}\|_{2}}{\sqrt{n}},\frac{\eta_{\zeta_{1}}\|\tilde{\bm{\mu}}_{ty}\|_{2}}{\sqrt{n}},\frac{\eta_{\zeta_{2}}}{\sqrt{n}}\|\tilde{X}\|_{n,2}\right)^{\top}.$
Following exactly the same analysis as Theorems 3 and 5, it can be proven that by choosing $\eta_{t},\eta_{ty},\eta_{\zeta_{1}}\asymp\sqrt{\frac{\log p}{n}},\eta_{\zeta_{2}}\asymp\sqrt{\frac{\log p\vee n}{n}}$, and $M_{r},M_{\pi},M_{\zeta}\asymp 1$ sufficiently large, $\hat{\tau}_{\textrm{DCal}}$ is an $\sqrt{n}$-consistent estimator for $\tau^{\ast}$ if either $s_{\zeta}=o(\sqrt{n}/\sqrt{\log p\log p\vee n})$ or $s_{\pi}\vee s_{r}=o(\sqrt{n}/\log p)$.